OpenStudy (anonymous):

list the common multiples of 2 and 3 up to 28

OpenStudy (anonymous):

fiddle!! D:

OpenStudy (anonymous):

Do you mean this ? 2*3=6 2^2*3=12 2^3*3=24 2*3^2=18

OpenStudy (anonymous):

I'm not exactly sure what they mean here.

OpenStudy (anonymous):

I think I did it right. http://www.mathsteacher.com.au/year7/ch03_prime/02_comm/comm.htm